# TrackPGD: A White-box Attack using Binary Masks against Robust Transformer Trackers

### v1

@@ 4 Experiments
We aim to examine TrackPGD’s impact on transformer tracker performance compared to other attack methods. We will start by outlining our experimental setup and protocols. Next, we will evaluate TrackPGD’s effectiveness on transformer trackers using object masks. The ablation study will illustrate the significance of our proposed difference loss for tracking and the impact of TrackPGD loss terms on overall performance. Additionally, we will showcase evaluation results based on predicted bounding boxes, highlighting TrackPGD’s advantages.
Evaluation setup
TrackPGD is compared with other existing attacks on VOT-STS2022 [12], VOT2018 [14] and VOT2016 [13] datasets. We conducted two primary experiments containing baseline, and unsupervised, on VOT-STS2022. In the baseline experiments, the Expected Overlap Analysis (EOA) and Accuracy-Robustness analysis are conducted using the anchor-based short-term tracking protocol. The unsupervised experiment computes the average accuracy in one initialization setting per sequence. The VOT2018 [14] and VOT2016 [13] are annotated with object bounding boxes and used here to evaluate the tracker’s robustness in bounding box prediction after applying the attacks. We selected three popular transformer-based trackers, MixFormerM [9], OSTrackSTS [10] and TransT-SEG [11], to assess their adversarial robustness against adversarial attacks. We applied four attack methods including CSA [4], SPARK [1], RTAA [3], and IoU [2], against these trackers. The TrackPGD perturbs the search region of the segmentation network in MixFormerM [9] and OSTrackSTS [10] while the search region of the tracking network remains untouched. The hyperparameters $(\lambda_{1},\lambda_{2})$ of the TrackPGD loss are determined through the validation experiment on the DAVIS2016 dataset [25], as explained in the Supplementary Materials.
@@ 4.1 Binary mask assessment
We evaluated the TrackPGD attack compared to other methods on two transformer trackers – MixFormerM [9] and OSTrackSTS [10] – using the VOT-STS2022 dataset, see  Table 2. Two streams of experiments from VOT-STS2022 (baseline and unsupervised) are conducted to indicate our attack performance under various protocols and metrics. Our findings demonstrate that TrackPGD was the most effective attack, reducing the robustness of MixFormerM [9] by around $91\%$ in the baseline experiment of VOT-STS2022 evaluation. TrackPGD also caused higher degradations to other evaluation metrics – EAO, accuracy, and AUC – than the two attack methods (i.e., IoU [2] and CSA [4] methods) against MixFormerM [9], both the on baseline and unsupervised streams of experiments. Among all measured metrics, TrackPGD ranked first by a significant margin on the VOT-STS2022 dataset. In attacking the OSTrackSTS tracker, TrackPGD also ranked first as an applicable attack per EAO, accuracy and AUC metrics. However, the IoU method [2] performs better than our method on the robustness metric. Figure 2 represents the performance of TrackPGD in generating inaccurate binary masks for MixFormerM [9] and OSTrackSTS [10]. The green color represents the original trackers’ output, while the red color represents the tracker output after the TrackPGD attack.
@@
Attack EAO Acc. Robust. AUC EAO Acc. Robust. AUC
@@
Table 2: The MixFormerM [9] and OSTrackSTS [10] results in generating object binary masks, when attacked by different methods on the VOT-STS2022 [12] dataset. Bold/italic are the first/second ranks, respectively.
Figure 2: The results of MixFormerM [9] and OSTrackSTS [10] attacked by TrackPGD. The green mask is the original output, while the red mask is the output after attack.
4.2 Ablation study
@@
Inspired by the SegPGD attack [7], TrackPGD relies on the predicted binary mask to generate the adversarial frames. The main challenge in following SegPGD for binary mask impairment to attack object trackers is that the SegPGD works well as a multi-class segmentation attack; while here, we have only two classes to switch the pixel labels. The object class and the background class pixels should switch their labels after TrackPGD. However, the object pixels are limited in comparison to the background pixels in almost all of the video frames. For this specific reason, it is necessary to define our method based on both object and background masks to involve all of the frame pixels in generating the adversarial perturbation. Plus, given the availability of two classes in the tracking scenario, we customized the segmentation loss using the tracking loss of the TransT [26] to effectively attack binary masks. We also demonstrate the role of ${L}_{\Delta}$ in TrackPGD versus using the vanilla SegPGD losses for the object and background masks, in Figure 3. The after attack masks show the essential role of our proposed combination loss $L_{\Delta}$ in crashing the object mask inferred by MixFormerM [9]. Although, the vanilla SegPGD losses also generate inaccurate binary masks, the mask impairment caused by $L_{\Delta}$ is significantly greater.
@@
Figure 3: The before (green) and after attack masks (red) generated by MixFormerM [9], while the vanilla SegPGDs and difference loss are used instead of binary cross entropy in the focal loss for the TrackPGD attack.
In addition, we conducted an ablation study to demonstrate the difference loss role in the focal loss. We predicted the binary masks by the MixFormerM tracker attacked by TrackPGD where the binary cross entropy is substituted with the vanilla SegPGD loss for the object mask, vanilla SegPGD for the background mask, and the difference loss $L_{\Delta}$. The DAVIS2016 [25] dataset was used to compute the percentage of the mean Jaccard index $J(M)$ of the binary masks indicating the average overlap between predicted and annotated binary masks. TrackPGD attack is stronger with the proposed difference loss, see Table 3, in comparison with the vanilla SegPGD losses for the object mask $G_{\tau}$, and the background mask $1-G_{\tau}$. The difference loss presented the most effective attack based on the generated binary masks by MixFormerM [9], which has the smaller $J(M)\%$ on average.
@@
$J(M)\%$ 85.82 52.86 37.28 30.30
Table 3: Cross-entropy substitutions in focal loss. The vanilla SegPGD losses versus the difference loss in the TrackPGD against MixFormerM [9] on DAVIS2016 [27].
@@
In this study, we show the effect of each loss term, dice and focal losses, in the TrackPGD attack. We conducted experiments on DAVIS2016 [25] using two evaluation metrics: the mean Jaccard index $J(M)$ and contour accuracy $F(M)$ of the predicted masks. The results were presented in Table 4 where the average $J\&F(M)$ percentage was used to demonstrate the impact of each losses term, focal vs. dice loss. Table 4 shows that the focal loss is more effective than the dice loss, resulting in smaller $J\&F(M)$ values. The $J\&F(M)$ value is used to measure the effectiveness of the attack, where a smaller value represents a more significant influence. We selected powers of 10 as coefficients of focal loss and powers of 2 as coefficients of dice loss. As in the literature, the loss related to IoU is normally 1 or 2 [11, 9], while the loss terms related to binary cross entropy (or focal loss) are selected as 1000 [20] or 1 [11]. An interesting observation in  Table 4 is that the MixFormerM is significantly impacted by TrackPGD, more than the TransT-SEG tracker using the same set of experiments. In both tables, the MixFormerM tracker achieves the smallest values of $J\&F(M)$ meaning more significant attack results.
$\lambda_{1}$ (with $\lambda_{2}=0$) $10000$ $1000$ $100$ $10$ $1$ $0.1$
@@
Table 4: $J\&F(M)\%$ values with varying focal and dice losses. Present the role of focal loss (changing $\lambda_{1}$ when $\lambda_{2}=0$) and dice loss (changing $\lambda_{2}$ when $\lambda_{1}=0$) in the TrackPGD attack. $J\&F(M)\%$ values are computed on DAVIS2016 [27] dataset.
4.3 Bounding box assessment
In  Table 5, we compare the performance of TrackPGD with other methods against the TransT-SEG [11] tracker. This evaluation is conducted on the VOT2018 [14] and VOT2016 [13] datasets, both of which contain annotated object bounding boxes in their video sequences. Higher values in the table indicate more frequent tracker failures after each attack, suggesting a more effective attack performance. Additionally, the average number of failures on the VOT2016 dataset illustrates TrackPGD’s capability in misleading TransT-SEG for object bounding box prediction. The average number of failures is calculated based on the failures on multiple evaluation runs. Despite being primarily designed to generate perturbations based on the network’s mask generation ability, TrackPGD introduces influential perturbations that rank it second in terms of robustness on both the VOT2016 and VOT2018 datasets. RTAA [3], the top-ranked method, is tailored to bounding box degradation using 1024 adversarial bounding boxes, while TrackPGD, aimed at flipping binary mask pixels, still emerges as the second most potent attack for bounding box degradation.
Attack Robustness (VOT2018) Robustness (VOT2016) Ave. failures (VOT2016)
No Attack 1.08 0.54 37.00
IoU [2] 2.14 1.87 115.47
CSA [4] 1.15 0.62 42.53
SPARK [1] 6.60 7.14 283.00
RTAA [3] 24.44 18.63 832.93
TrackPGD 10.76 10.02 483.00
Table 5: The TransT-SEG [11] robustness in predicting object bounding box, attacked by different methods on the VOT2018 [14] and VOT2016 [13] datasets.
We also illustrate the bounding box evaluation of TransT-SEG tracker [11] in several subsequences from VOT2018 [14] in Figure 4. The original annotation is depicted in green, representing a rotated bounding box per frame. The deviation from the actual performance of the TransT-SEG tracker (dark blue box) is comparable to the other attack approaches. The TrackPGD method is indicated in white, which shows a rapid and significant deviation even from the initial frames of the video sequence.

### v2

@@ 5 Experiments
Datasets.
Several video segmentation and tracking datasets. including DAVIS 2016 (Perazzi et al., 2016), VOT2022STS (Kristan et al., 2023), VOT2018 (Kristan et al., 2018), GOT10k (Huang et al., 2019), and UAV123 (Mueller et al., 2016), are adopted in our experiments.
Victim Trackers.
We selected three popular transformer-based trackers, MixFormerM (Cui et al., 2022), OSTrackSTS (Ye et al., 2022), TransT-SEG (Chen et al., 2023) and one non-transformer tracker, RTS (Paul et al., 2022), to assess their adversarial robustness against adversarial attacks.
Adversarial Attacks.
We applied four attack methods including CSA (Yan et al., 2020), SPARK (Guo et al., 2020), RTAA (Jia et al., 2020), and IoU (Jia et al., 2021), against these trackers. The TrackPGD perturbs the search region of the segmentation network in MixFormerM (Cui et al., 2022) and OSTrackSTS (Ye et al., 2022) while the search region of the tracking network remains untouched. The hyperparameters $(\lambda_{1},\lambda_{2})$ of the TrackPGD loss are determined through the validation experiment on the DAVIS2016 (Perazzi et al., 2016), as explained in the section Fine tuning the parameters.
Metrics.
The expected average overlap (denoted as EAO and AO), robustness, accuracy and AUC from VOT2022 metrics, Mean Jaccared Index $J(M)\%$ from DAVIS2016 evaluation toolkit, AUC and precision rates for GOT01k and UAV123 datasets are used to rank the attack performance.
@@ 5.1 Binary mask assessment
Quantitative evaluation.
We evaluated the TrackPGD attack compared to other methods on two transformer trackers – MixFormerM (Cui et al., 2022) and OSTrackSTS (Ye et al., 2022) – using the VOT-STS2022 dataset, see Table 2. Two streams of experiments from VOT-STS2022 (baseline and unsupervised) are conducted to indicate our attack performance under various protocols and metrics. Our findings demonstrate that TrackPGD was the most effective attack, reducing the robustness of MixFormerM (Cui et al., 2022) by around $91\%$ in the baseline experiment of VOT-STS2022 evaluation. TrackPGD also caused higher degradations to other evaluation metrics – EAO, accuracy, and AUC – than the two attack methods (i.e., IoU (Jia et al., 2021) and CSA (Yan et al., 2020) methods) against MixFormerM (Cui et al., 2022), both on baseline and unsupervised streams of experiments. Among all measured metrics, TrackPGD ranked first by a significant margin on the VOT-STS2022 dataset. In attacking the OSTrackSTS tracker, TrackPGD also ranked first as an applicable attack per EAO, accuracy and AUC metrics. However, the IoU method (Jia et al., 2021) performs better than our method on the robustness metric. To check further examples of mask prediction after attacks, we refer the reader to the Supplementary materials.
Table 2: The MixFormerM (Cui et al., 2022), OSTrackSTS (Ye et al., 2022) results in generating object binary masks, when attacked by different methods on the VOT-STS2022 (Kristan et al., 2023) dataset. TrackPGD is the only applicable white-box attack on MixFormerM and OSTrackSTS trackers. Bold/italic are the first/second ranks, respectively. The other white-box attacks, RTAA and SPARK are not applicable to these trackers due to difference in backbones. For further information, please check the text.
@@
Tracker Attack EAO Accuracy Robust. AUC EAO Accuracy Robust. AUC
@@
5.2 Bounding box assessment
Quantitative evaluation.
Table 5 presents the attack results on the TransT-SEG (Chen et al., 2023) and RTS (Paul et al., 2022) trackers using the GOT-10k dataset (Huang et al., 2019). Although our method focuses on manipulating object binary masks, the reported numbers for bounding box evaluation are very close to those of competitive attacks aimed at misleading trackers into providing inaccurate bounding boxes. We surpassed the RTAA attack on the GOT-10k dataset in both success rate scores. However, RTAA performs better in terms of average overlap (AO). TrackPGD scores are also very close to SPARK scores, the most competitive attack using object bounding boxes on GOT-10k (Huang et al., 2019), in misleading trackers to predict accurate object bounding boxes. Despite our method’s aim to misclassify object and background pixels, it proves to be a highly competitive approach compared to bounding box-centered attacks. Furthermore, our TrackPGD has an extra advantage over other white-box attacks which is that other attack opponents, RTAA and SPARK, are not applicable as a white-box attack on new transformer and non-transformer trackers containing MixFormerM, OSTrackSTS and RTS. For the RTS tracker (Bhat et al., 2020), though, TrackPGD is the only white-box attack which is applicable and significantly affect the tracker performance.
The TransT-SEG (Chen et al., 2023) and RTS (Paul et al., 2022) trackers robustness after a white-box attack is also assessed on the UAV123 dataset in Table 8. Once again, our method achieves scores that are closely competitive with SPARK (Guo et al., 2020) and RTAA (Jia et al., 2020) attacks on bounding box prediction, even though TrackPGD is developed to mislead trackers in predicting binary masks. Also, our method is the first white-box attack applicable on RTS tracker and presents a very effective performance in decreasing the RTS scores on the UAV123 (Mueller et al., 2016) dataset. To check further examples of bounding box prediction after attacks, we refer the reader to the Supplementary materials.
Table 3: TransT-SEG (Chen et al., 2023) and RTS (Paul et al., 2022) trackers evaluation after white-box applicable approaches, SPARK (Guo et al., 2020), RTAA (Jia et al., 2020) and TrackPGD (ours), on the GOT10k (Huang et al., 2019) dataset.
Tracker Attacker AO SR0.5 SR0.75
TransTSEG No Attack 0.719 0.816 0.680
RTAA 0.127 0.115 0.060
SPARK 0.120 0.068 0.015
TrackPGD (ours) 0.139 0.092 0.025
RTS No Attack 0.736 0.830 0.681
TrackPGD (ours) 0.292 0.291 0.162
Table 4: Evaluation of our proposed TrackPGD vs. other white-box attacks applied on TransT-SEG (Chen et al., 2023) and RTS (Paul et al., 2022) trackers on the UAV123 dataset.
Tracker Attacker AUC Precision
TranstSEG No attack 0.670 0.859
SPARK 0.050 0.074
RTAA 0.058 0.077
TrackPGD 0.135 0.205
RTS No attack 0.666 0.873
TrackPGD 0.427 0.674
6 Conclusion
Existing adversarial attacks for object trackers are ineffective as white-box attacks against transformer trackers. To evaluate the adversarial robustness of transformer trackers, new white-box attacks that utilize the tracker’s own gradients are necessary. This paper introduces TrackPGD, a novel white-box attack that leverages object binary masks to assess the adversarial robustness of transformer trackers. We highlight the effectiveness of our proposed difference loss in impacting tracker performance compared to standard segmentation losses such as SegPGD. The efficacy of TrackPGD is validated through comprehensive experiments on various transformer and non-transformer networks on popular datasets. Despite the strong performance of our method, TrackPGD is limited to object trackers that generate a binary mask. Future work could extend these attacks to a broader range of trackers. Additionally, exploring defense strategies, such as adversarial purification with generative models (e.g., GANs or diffusion models) or adversarial training, could improve adversarial robustness in object tracking networks.
Acknowledgments
Appendix A Supplementary materials
A.1 Algorithm
The principle steps of TrackPGD algorithm is presented in Algorithm 1. The sign of step size $\alpha$, line 9 in Algorithm 1, is determined according the best performance of each tracker, following the decoupling of the norm and direction of gradients in adversarial attacks (Rony et al., 2019).
2:${I}_{\text{adv}}^{0}\leftarrow I_{\tau}$triangleright initialization
4:for $t=1\ldots T$ do
6:     $L_{\Delta}\leftarrow L_{\text{SegPGD}}(M^{t},G_{\tau})-L_{\text{SegPGD}}(M^{t},1-G_{\tau})$triangleright compute difference of SegPGD losses
7:     $L_{\text{focal}}\leftarrow\alpha_{t}(1-p_{t})^{\gamma}L_{\Delta}$triangleright compute focal loss
8:     $L_{\text{dice}}\leftarrow 1-2\,\text{IoU}(M^{t},G_{\tau})$triangleright compute dice loss
10:     ${I}_{\text{adv}}^{t}\leftarrow I_{\text{adv}}^{t-1}+\alpha\,\mathrm{sign}(\nabla_{I_{\text{adv}}^{t-1}}L)$triangleright update adversarial example
12:end for
A.2 Ablation study
@@
Inspired by the SegPGD attack (Gu et al., 2022), TrackPGD relies on the predicted binary mask to generate the adversarial frames. The main challenge in following SegPGD for binary mask impairment to attack object trackers is that the SegPGD works well as a multi-class segmentation attack; while here, we have only two classes to switch the pixel labels. The object class and the background class pixels should switch their labels after TrackPGD. However, the object pixels are limited in comparison to the background pixels in almost all of the video frames. For this specific reason, it is necessary to define our method based on both object and background masks to involve all of the frame pixels in generating the adversarial perturbation. Plus, given the availability of two classes in the tracking scenario, we customized the segmentation loss using the tracking loss of the TransT (Chen et al., 2021) to effectively attack binary masks. We also demonstrate the role of ${L}_{\Delta}$ in TrackPGD versus using the vanilla SegPGD losses for the object and background masks, in Figure 2. The after attack masks show the essential role of our proposed combination loss $L_{\Delta}$ in crashing the object mask inferred by MixFormerM (Cui et al., 2022). Although, the vanilla SegPGD losses also generate inaccurate binary masks, the mask impairment caused by $L_{\Delta}$ is significantly greater.
Figure 2: The before (green) and after attack masks (red) generated by MixFormerM (Cui et al., 2022), while the vanilla SegPGDs and difference loss are used instead of binary cross entropy in the focal loss for the TrackPGD attack.
@@
Table 5: The impact of cross-entropy substitutions in focal loss on $J(M)\%$. The vanilla SegPGD losses versus the difference loss in the TrackPGD against MixFormerM (Cui et al., 2022) on DAVIS2016 (Mueller et al., 2016).
@@
85.82 52.86 37.28 30.30
In addition, we conducted an ablation study to demonstrate the difference loss role in the focal loss. We predicted the binary masks by the MixFormerM tracker attacked by TrackPGD where the binary cross entropy is substituted with the vanilla SegPGD loss for the object mask, vanilla SegPGD for the background mask, and the difference loss $L_{\Delta}$. The DAVIS2016 (Perazzi et al., 2016) dataset was used to compute the percentage of the mean Jaccard index $J(M)$ of the binary masks indicating the average overlap between predicted and annotated binary masks. TrackPGD attack is stronger with the proposed difference loss, see Table 5, in comparison with the vanilla SegPGD losses for the object mask $G_{\tau}$, and the background mask $1-G_{\tau}$. The difference loss presented the most effective attack based on the generated binary masks by MixFormerM (Cui et al., 2022), which has the smaller $J(M)\%$ on average.
@@
In this study, we show the effect of each loss term, dice and focal losses, in the TrackPGD attack. We conducted experiments on DAVIS2016 (Perazzi et al., 2016) using two evaluation metrics: the mean Jaccard index $J(M)$ and contour accuracy $F(M)$ of the predicted masks. The results were presented in Table 6 where the average $J\&F(M)$ percentage was used to demonstrate the impact of each losses term, focal vs. dice loss. Table 6 shows that the focal loss is more effective than the dice loss, resulting in smaller $J\&F(M)$ values. The $J\&F(M)$ value is used to measure the effectiveness of the attack, where a smaller value represents a more significant influence. We selected powers of 10 as coefficients of focal loss and powers of 2 as coefficients of dice loss. As in the literature, the loss related to IoU is normally 1 or 2 (Chen et al., 2023; Cui et al., 2022), while the loss terms related to binary cross entropy (or focal loss) are selected as 1000 (Yan et al., 2021) or 1 (Chen et al., 2023). An interesting observation in Table 6 is that the MixFormerM is significantly impacted by TrackPGD, more than the TransT-SEG tracker using the same set of experiments. In both tables, the MixFormerM tracker achieves the smallest values of $J\&F(M)$ meaning more significant attack results.
Table 6: $J\&F(M)\%$ values with varying focal and dice losses. Present the role of focal loss (changing $\lambda_{1}$ when $\lambda_{2}=0$) and dice loss (changing $\lambda_{2}$ when $\lambda_{1}=0$) in the TrackPGD attack. $J\&F(M)\%$ values are computed on DAVIS2016 (Mueller et al., 2016) dataset.
$\lambda_{1}$ ($\lambda_{2}=0$) $10000$ $1000$ $100$ $10$ $1$ $0.1$
@@
A.3 Fine tuning the parameters
We used the DAVIS2016 dataset as the validation set to adjust the $\lambda_{1}$ and $\lambda_{2}$ parameters. The DAVIS protocol calculates the Jaccard Index $\mathcal{J}$ and contour accuracy $\mathcal{F}$ via three error statistic measures: mean, recall, and decay. The Jaccard index measures the Intersection over Union (IoU) metric, while the contour accuracy is computed by calculating the F-measure of the predicted mask and the ground truth. To obtain the F-measure, the predicted mask is considered as a set of closed contours, and precision $P_{c}$ and recall $R_{c}$ are calculated between contour points by comparing these contours with the ground-truth contours. Then, the F-measure $\frac{2P_{c}R_{c}}{P_{c}+R_{c}}$ is computed to represent the accuracy of the predicted mask. To fine-tune the TrackPGD parameters, we computed the average value of $\mathcal{J}(R)$ and $\mathcal{F}(R)$, represented as $\mathcal{J}\&\mathcal{F}(R)$. This approach considers both IoU and contour accuracy over time across video frames. Since this measurement is close to the robustness definition in VOT challenge (Kristan et al., 2023), we chose this metric to identify the best hyperparameters to determine the optimal coefficient of TrackPGD loss.
The following heatmaps show the average values of the Jaccard index $\mathcal{J}$ and contour accuracy $\mathcal{F}$ for the tracker’s recall metric. The better performance of the tracker is represented by the greater $\mathcal{J}\&\mathcal{F}(R)$. As our reported numbers are based on attack results, a smaller number indicates a more effective attack. Therefore, the best parameters for MixFormerM are $(\lambda_{1},\lambda_{2})=(10,2.5)$, while the TrackPGD is strongest against OSTrackSTS with $(\lambda_{1},\lambda_{2})=(1,10)$, against TransT-SEG with $(\lambda_{1},\lambda_{2})=(50,2.5)$ and against RTS with $(\lambda_{1},\lambda_{2})=(1,5)$. It is also worth mentioning that optimal performance is achieved when both terms are active at intermediate values, as shown in Figure 3. This suggests that neither term alone yields the best results.
Figure 3: The TrackPGD parameter heatmaps from left to right, first row: MixFormerM(Cui et al., 2022; Yan et al., 2021), OSTrackSTS (Ye et al., 2022; Yan et al., 2021) and second row: TransT-SEG (Chen et al., 2023), RTS (Paul et al., 2022) trackers.
A.4 Binary mask assessment
Qualitative evaluation.
Figure 4 represents the performance of TrackPGD in generating inaccurate binary masks for MixFormerM (Cui et al., 2022) and OSTrackSTS (Ye et al., 2022) and RTS (Paul et al., 2022) trackers on the successive frames of several video sequences. The green color represents the original trackers’ output, while the red color represents the tracker output after the TrackPGD attack. As it is presented in Table 2, the effect of the TrackPGD attack on the MixFormerM performance is more significant.
Figure 4: The results of MixFormerM (Cui et al., 2022), RTS (Paul et al., 2022), and OSTrackSTS (Ye et al., 2022) attacked by TrackPGD. The green mask is the original output, while the red mask is the after attack mask.
A.5 Bounding box assessment
Quantitative evaluation.
Table 5 presents the attack results on the TransT-SEG (Chen et al., 2023) and RTS (Paul et al., 2022) trackers using the GOT-10k dataset (Huang et al., 2019). Although our method focuses on manipulating object binary masks, the reported numbers for bounding box evaluation are very close to those of competitive attacks aimed at misleading trackers into providing inaccurate bounding boxes. We surpassed the RTAA attack on the GOT-10k dataset in both success rate scores. However, RTAA performs better in terms of average overlap (AO). TrackPGD scores are also very close to SPARK scores, the most competitive attack using object bounding boxes on GOT-10k (Huang et al., 2019), in misleading trackers to predict accurate object bounding boxes. Despite our method’s aim to misclassify object and background pixels, it proves to be a highly competitive approach compared to bounding box-centered attacks. Furthermore, our TrackPGD has an extra advantage over other white-box attacks which is that other attack opponents, RTAA and SPARK, are not applicable as a white-box attack on new transformer trackers containing MixFormerM, OSTrackSTS and RTS. For the RTS tracker (Bhat et al., 2020), though, TrackPGD is the only white-box attack which is applicable and significantly affect the tracker performance.
The TransT-SEG (Chen et al., 2023) and RTS (Paul et al., 2022) trackers robustness after a white-box attack is also assessed on the UAV123 dataset in Table 8. Once again, our method achieves scores that are closely competitive with SPARK (Guo et al., 2020) and RTAA (Jia et al., 2020) attacks on bounding box prediction, even though TrackPGD is developed to mislead trackers in predicting binary masks. Also, our method is the first white-box attack applicable on RTS tracker and presents a very effective performance in decreasing the RTS scores on the UAV123 (Mueller et al., 2016) dataset.
Table 7: TransT-SEG (Chen et al., 2023) and RTS (Paul et al., 2022) trackers evaluation after white-box applicable approaches, SPARK (Guo et al., 2020), RTAA (Jia et al., 2020) and TrackPGD (ours), on the GOT10k (Huang et al., 2019) dataset.
Tracker Attacker AO SR0.5 SR0.75
TransTSEG No Attack 0.719 0.816 0.680
RTAA 0.127 0.115 0.060
SPARK 0.120 0.068 0.015
TrackPGD (ours) 0.139 0.092 0.025
RTS No Attack 0.736 0.830 0.681
TrackPGD (ours) 0.292 0.291 0.162
Table 8: Evaluation of our proposed TrackPGD vs. other white-box attacks applied on TransT-SEG (Chen et al., 2023) and RTS (Paul et al., 2022) trackers on the UAV123 dataset.
Tracker Attacker AUC Precision
TranstSEG No attack 0.670 0.859
SPARK 0.050 0.074
RTAA 0.058 0.077
TrackPGD 0.135 0.205
RTS No attack 0.666 0.873
TrackPGD 0.427 0.674
Qualitative evaluation.
We also illustrate the bounding box evaluation of TransT-SEG tracker (Chen et al., 2023) in several frames from the VOT2018 (Kristan et al., 2018) sequence in Figure 5. The greed/red colors indicates the before and after attack predictions. Our proposed method, TrackPGD, results in misleading the tracker while its perturbation is generated only based on the binary mask. Other attacks such as SPARK and RTAA, which fully built upon bounding box(es) also significantly degrade the target bounding boxes. As the quantitative results in Table 7 indicated, the IoU (Jia et al., 2021) and CSA (Yan et al., 2020) attack, black-box attacks, are behind the white-box methods’ performance.